a manufacturer plans to spend $150,000 on research and development for a new lawn mower and then $200 to manufacture each mower. the formula C=200n+150,000 gives the cost in dollars of n mowers. what is the cost of 5000 mowers?
what is the cost of 5001 mowers?
by how much did the one extra mower increase the cost?

Answer by nyc_function(2741)

C = 200n + 150,000
Let n = 5000
C = 200(5000) + 150,000
C = $1,150,000
We now do the same thing for n = 5001 and solve for C.
C = 200(5001) + 150,000
C = $1,150,200
Let E = money that extra mover increased the cost
E = 1,150,200 - 1,150,000
E = $1,000,200
I hope this helps.